OpenStudy (anonymous):

The width of a rectangle is 5ft. less than the length. The area of the rectangle is 300 ft ^2. Find the length and width of the rectangle.

OpenStudy (anonymous):

w = L - 5 (first sentence) L*w = 300 (formula for area of a rectangle) substitute in first sentence says what w = so put that in the second L(L - 5) = 300 L^2 - 5L = 300 L^2 - 5L - 300 = 0 (L - 20)(L + 15) = 0 L = 20 or L = -15 Because Length must be positive.. L = 20 W = L - 5 W = 20 - 5 W = 15